If 500 g of carbon disulfide burns in the presence of oxygen to produce carbon dioxide and sulfur dioxide, how many liters of sulfur dioxide collected over water measured at 27°C and 740-mmHg, are produced?
1
Expert's answer
2020-03-30T10:53:57-0400

"CS_2 + 3O_2 \rightarrow CO_2 + 2SO_2"


Find moles of "SO_2"


"500 g CS_2 (\frac{1 mole CS_2}{76.14 g CS2})(\frac{2 moleSO_2}{1 mole CS_2}) = 13.1 mol SO_2"


Use Ideal Gas Law to find the volume of "SO_2"

"PV = nRT"

"V =\frac{nRT}{P}"


"n=13.1 mol"

"T(K) = 273 + 27 = 300 K"

"P(atm) = 740 mm Hg (\frac{1 atm}{760 mm Hg})= 0.974 atm"

"R = 0.0821 \frac{L\times atm}{mol\times K}"


"V = \frac{13.1\times 0.0821 \times 300}{0.974} = 331 L"